40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2363     Entry time: Tue Dec 8 03:53:49 2009
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: Free swinging spectra of ETMX  

In this night, I checked the free swinging spectra of ETMX to make sure nothing wrong with ETMX by the wiping.

Compared with the past (Aug.6 2008), the spectra of ETMX doesn't show significant change.

Successfully the wiping activity didn't change its configuration so much and didn't bring bad situations.

(bad situation means for example, the suspended components hit some others).

 

 The spectra of ETMX by DTT are attached. Also you can see the past spectra in Yoichi's entry.

Yoichi's data was taken during the air-pressure condition, so it's good for comparing.

Actually I compared those data by my eyes, because I could not get the past raw data somehow.

The resonant frequencies and their typical height changed a little bit, but I think those are not significant.

NOTE: In the figure, pitch and yaw modes (~0.57Hz and ~0.58Hz) look like having a smaller Q-factor than the past.

 

Attachment 1: ETMX_air.png  365 kB  | Hide | Hide all
ETMX_air.png
ELOG V3.1.3-